If f(x)= x^2+bx+c x=2 is one of the roots of f(x)=0 so f(2) =
1
Expert's answer
2019-09-10T13:22:10-0400

Let us consider a function "f(x)=x^2+bx+c".

Solutions of the equation "f(x)=0" satisfy the Vieta's formulas


"x_1+x_2=-b, \quad x_1x_2=c"

In our case


"2+x_2=-b, \quad 2x_2=c"

From these equations we get


"2b+c=-4"

Finally


"f(2)=2^2+2b+c=4+2b+c=4-4=0"

"f(2)=0"